I guess if you have a frequency counter with a GPIB interface or a simple flip-flop XOR based phase/frequency discriminator, you can feed the output to a 3113 and use at as an input to a slow EPICS PID to bring the beat frequency to within range.
Actually, we need a frequency discriminator for the Green Locking so it might be good to brainstorm about this with Aidan and Koji.
good idea - we now know where we have to be +- acouple of 100mK as the impact of changes in room temp is not equal for both systems so we have to slightly adjust the temperature of one cavity.
Right now we have a much larger problem with the AOM. if we lock both cavities and the AOM frequency changes slowly over time (both cavities drift a bit) we get a huge amplitude modulation for the refcav, meaning a drift in power from almost 0 to 2.5V (the full range) topped by a sine-like modulation, looks like an etalon with about 10% modulation depth, every 1.4Mhz. It's not the power of the VCO, that changes only by about .2%, but you see this modulation there as well, but almost covered by the noise of the DAQ already. It's a power modulation of the diffracted beam as you can see it in the reflected and transmitted light with the same sign, so it's not pointing or so. I think it could be an impedance matching problem causing some standing wave resonator. We try to investigate it today. I will post a graph later today...